# Large Language Model-Assisted UAV Operations and Communications: A Multifaceted Survey and Tutorial

### v1

@@ II-F1 Edge, On-Device, and Hybrid LLM Deployment
    Centralized Edge Inference: In this approach, UAVs transmit sensor data to an edge server or ground station hosting the LLM for inference. Although this paradigm is communication-intensive, techniques such as token reduction can help mitigate transmission overhead. Deploying LLMs at the edge typically requires model compression, quantization, pruning, and parameter-efficient fine-tuning to accommodate constrained hardware resources. When properly optimized, edge-based inference can support real-time decision-making, which is essential for safety-critical UAV operations.
@@
    Local Inference (On-Device): Executing LLMs directly on UAV hardware improves data privacy, minimizes latency, and increases robustness against intermittent connectivity. Prior work, such as Aero-LLM [22], has demonstrated the feasibility of deploying very small models (e.g., OPT-125M/350M) on UAVs. However, such lightweight models often lack the reasoning depth required for complex tasks, including trajectory planning and resource allocation based on ground user requirements. On-device LLM deployment is further constrained by three key factors: (i) high computational demands, where generating even a single token may take several seconds on embedded platforms; (ii) excessive energy consumption, which can rapidly deplete UAV batteries; and (iii) stringent storage and memory limitations, as model parameters and intermediate activations frequently exceed the Random Access Memory (RAM) capacity of typical UAV systems [70].
@@
    Split Inference (Model Partitioning): In split inference, the LLM is partitioned such that the initial layers execute onboard the UAV to process raw sensor data. In contrast, deeper layers run on edge servers or cloud platforms. This strategy substantially reduces onboard computational requirements and limits data transmission to compact intermediate representations, rather than high-resolution raw sensor inputs.
@@
    Collaborative Device–Server Inference: This paradigm leverages cooperation between lightweight onboard models and powerful server-side LLMs. Techniques such as speculative decoding enable a small local model to generate preliminary outputs, which are subsequently verified or refined by a larger remote model. Compared to transmitting raw multimodal data, this approach significantly reduces bandwidth requirements and is particularly well suited to environments with limited or unreliable communication infrastructure [93].
While centralized edge inference can enable real-time LLM processing for UAV applications, its overall effectiveness is fundamentally constrained by the memory capacity, computational resources, and energy availability of edge hardware. These limitations directly influence the scalability and practicality of deploying LLMs in UAV systems. Table V compares LLM inference paradigms for UAVs.
II-F2 Balancing Model Size, Memory, and Performance on the Edge
The scalability of LLMs on edge hardware is fundamentally constrained by memory capacity, computational throughput, and energy availability, all of which are significantly more limited than in cloud environments. A theoretical upper bound on the number of model parameters (P) that can be accommodated on an edge device can be approximated as $P=\frac{M\times 8}{b}$, where $M$ denotes the available memory in bytes and $b$ represents the bit precision per parameter. Under this formulation, an edge device with 8–16 GB of usable RAM can theoretically support approximately 4–8 billion parameters in FP16 precision, or 16–32 billion parameters when employing 4-bit quantization.
@@ II-F2 Balancing Model Size, Memory, and Performance on the Edge
In practice, however, this upper bound is substantially reduced by runtime overheads, activation storage, and intermediate buffers, lowering the adequate capacity to roughly 1–10 billion parameters. This practical range is consistent with the scale of current state-of-the-art on-device models, such as Gemini Nano [20] and Qualcomm’s LLMs with up to 10 billion parameters [120]. Consequently, most real-world edge deployments operate in the 1–3 billion-parameter regime, whereas larger models require cloud offloading or distributed execution across multiple devices [72, 36].
Given the stringent memory, computational, and energy constraints of edge platforms, achieving real-time LLM inference on UAVs typically requires not only careful model scaling and optimization, but also high-speed, low-latency communication with edge or cloud servers. 5G networks, particularly through their URLLC capabilities and high-throughput links, provide a critical enabling infrastructure. By enabling rapid computation offloading and seamless access to larger models, 5G connectivity enables UAVs to maintain responsive, context-aware decision-making even in highly dynamic, resource-constrained environments.
Table VI: Techniques for scaling LLMs to extended context windows in UAV applications, comparing selective context filtering, prompt compression, memory mechanisms, and AGI-assisted designs for long-horizon mission planning and interaction.
Technique Purpose Key Mechanism Pros Notes
Selective Context Mitigate long-context processing limitations Filters redundant/low-information tokens using token-level self-information Reduces memory usage and computational overhead; preserves core semantic content; accelerates text generation Minimal performance degradation, but may discard some less obvious informative content
LLMLingua Long-context prompt compression with semantic preservation Hierarchical prompt compression: (i) budget controller assigns compression ratios; (ii) token-level iterative compressor models conditional dependencies Stronger compression than Selective Context; better semantic fidelity; retains critical information More complex; may require additional preprocessing steps; computational overhead during compression
Short-Term Working Memory Maintain local coherence in extended interactions Dynamic buffer for recent observations, actions, intermediate states Ensures short-term continuity; improves immediate reasoning Limited to recent context; cannot store long-term mission knowledge
Long-Term Summary Memory Preserve extended knowledge over long missions Compact, abstracted representation of key patterns, insights, accumulated knowledge Supports strategic reasoning; reduces contextual drift over long durations Abstracted nature may lose fine-grained details
AGI-Assisted LLM Designs Enhance adaptive context retention beyond fixed windows Episodic and semantic memory mechanisms for dynamic storage, retrieval, and updating of information Improves consistency and coherence in long-horizon interactions; adaptive learning Complex architectures; higher computational and design requirements; research-level approach
II-F3 Scaling LLMs for Extended Contexts
LLMs struggle to process long contexts efficiently, which poses a fundamental limitation for real-world applications such as extended UAV missions and long-duration conversational interactions. This challenge stems from the architectural design of Transformer-based models, whose self-attention mechanism scales quadratically with the context length, leading to both computational complexity and memory consumption. Consequently, processing long inputs becomes prohibitively expensive, often necessitating aggressive truncation that can degrade task performance. This limitation is further reinforced during pre-training, as most LLMs are trained on fixed-length context windows, inherently constraining their ability to maintain broader or persistent contextual understanding. Selective Context is an effective strategy to mitigate this issue by filtering out redundant or low-information content before the LLM processes it. By estimating token-level self-information, this approach identifies and discards less informative tokens, thereby reducing memory usage and computational overhead while preserving the core semantic content. Empirical evaluations demonstrate that Selective Context can substantially reduce GPU memory consumption and accelerate text generation with only minimal performance degradation, enabling LLMs to process long documents and extended conversations without severe truncation [67].
A related and more structured approach, LLMLingua, addresses the long-context challenge through hierarchical prompt compression. LLMLingua operates in two stages: first, a budget controller assigns different compression ratios to various components of the prompt (e.g., task instructions, demonstrations) to preserve semantic fidelity; second, a token-level iterative compressor further reduces the prompt by modeling conditional dependencies to retain critical information. Compared to Selective Context, LLMLingua achieves stronger compression while better preserving semantic meaning [52].
Beyond prompt compression, architectural solutions can further alleviate long-context limitations by integrating explicit memory mechanisms. In particular, a conversation memory module can be composed of two complementary components: a) Short-Term Working Memory- A dynamic buffer that maintains recent observations, actions, and intermediate states, ensuring local coherence and continuity; b) Long-Term Summary Memory- A compact and abstracted representation that stores key patterns, strategic insights, and accumulated knowledge over extended mission durations. In addition, Artificial General Intelligence (AGI)-assisted LLM designs can further enhance context retention and adaptive learning capabilities. Unlike conventional LLMs constrained by fixed context windows, AGI-inspired models incorporate episodic and semantic memory mechanisms that dynamically store, retrieve, and update information over time. Such mechanisms reduce contextual drift and improve consistency and coherence in long-horizon interactions and decision-making processes [50]. Table VI compares techniques for scaling LLMs.
II-F4 Structuring Knowledge for Reliable Retrieval
@@
Chunking is a fundamental component of modern RAG systems. As LLM context windows continue to grow, the primary bottleneck has shifted from generation to retrieval, making the structure and granularity of chunks central to overall system performance. Effective chunking breaks information into coherent and meaningful units that reflect how humans naturally organize and understand ideas, rather than relying on arbitrary text boundaries. This distinction is critical because the retrieval quality determines the context the model ultimately sees. Most RAG hallucinations arise from missing, fragmented, or poorly segmented information.
Different chunking strategies, such as fixed-length, hierarchical, semantic, sliding-window, sentence-based, topic-boundary, and hybrid approaches, offer distinct strengths and tradeoffs, and their effectiveness depends heavily on the structure and nature of the underlying content. Ultimately, chunking is not merely a preprocessing step; it functions as an information architecture for AI, shaping how models perceive, retrieve, and reason over knowledge. Better chunking leads directly to more accurate, efficient, and reliable RAG systems.
II-F5 Quantization as an Inference Optimization Technique for Efficient LLMs
@@
Quantization is a widely used model compression technique that improves the efficiency of LLMs by reducing the numerical precision of their parameters, and, in some cases, their activations, from high-precision floating-point formats (e.g., FP32 or FP16) to lower-bit representations such as INT8, INT4, or INT3. This reduction in precision substantially decreases model size, memory footprint, computational complexity, inference latency, and energy consumption, making quantization particularly important for deploying LLMs on resource-constrained edge devices, including Raspberry Pi–class hardware. Quantization can be applied after training via Post-Training Quantization (PTQ), which converts pretrained model weights without retraining, or during training via Quantization-Aware Training (QAT), which simulates low-precision arithmetic to preserve accuracy better. Despite these efficiency gains, quantization introduces tradeoffs: aggressive reductions in bit-width may lead to accuracy degradation, numerical instability, or additional dequantization overhead, particularly for tasks involving complex reasoning or arithmetic. As a result, effective quantization requires carefully balancing efficiency gains with acceptable performance loss to enable practical, sustainable edge AI deployment.
II-F6 Distillation and Prunning
Distillation provides a practical and scalable approach to compressing LLMs by enabling smaller student models to learn from the outputs and internal representations of a powerful teacher, rather than relying solely on task-specific supervision. It can be tightly integrated with techniques such as low-rank adaptation and supernet training, thereby requiring only a small number of additional parameters to train while keeping the pretrained weights frozen. This strategy enables efficient derivation of multiple model variants with different sizes, latencies, and accuracy tradeoffs from a single training run. As a result, distillation not only reduces model size and training cost but also supports dynamic deployment across heterogeneous edge devices, making it particularly well-suited to enterprise and edge AI scenarios where efficiency, flexibility, and performance must be carefully balanced. In parallel, LLM pruning further streamlines models by removing redundant or less important components, producing smaller, faster networks while preserving performance.
@@ III LLM for UAV Communications
Due to their high mobility and flexible deployment, UAVs have been widely adopted in civil and commercial applications such as environmental monitoring, traffic management, parcel delivery, and precision agriculture [82, 99, 46, 68]. Beyond sensing and service roles, UAVs increasingly function as airborne communication relays for ground sensors in infrastructure-limited or emergency scenarios, including disaster response and border surveillance. By dynamically repositioning to establish short-range line-of-sight (LoS) links, UAVs enable reliable and high-throughput data transmission while extending network coverage beyond fixed terrestrial infrastructure. These capabilities make UAV-assisted networks particularly suitable for remote or hazardous environments, where rapid data collection and wide-area coverage are required with minimal human intervention [31].
LLMs are well suited to UAV communication systems because they directly address key challenges such as dynamic network topology, heterogeneous information, and time-sensitive decision-making. Unlike conventional optimization or reinforcement learning approaches that rely on precise modeling or extensive retraining, LLMs can reason over high-level task descriptions, historical solutions, and real-time feedback to generate communication-aware decisions with low iteration overhead. This section reviews recent efforts to integrate LLMs into UAV navigation, swarm coordination, safety, and network optimization, highlighting their emerging role in intelligent aerial communication systems.
@@ III-A LLM-Assisted UAV Navigation, Planning, and Placement
Addressing this gap, Yuan et al.[129] introduce Next-Generation LLM for UAV (NELV), the first end-to-end framework that connects natural-language instructions directly to fully executable UAV flight trajectories across short-, medium-, and long-range missions. NELV integrates five key components, LLM-as-Parser, route planning, path planning, control execution, and real-time monitoring, to enable complex tasks such as multi-UAV patrol, multi-point-of-interest delivery, and long-distance multi-hop relocation. Through three representative use cases, the framework demonstrates how LLMs can significantly reduce pilot workload while adhering to regulatory, airspace, and environmental constraints. Furthermore, the authors propose a five-level autonomy roadmap that outlines the progression from current language-based instruction parsing to fully autonomous LLM-as-Autopilot systems, identifying critical challenges related to regulation-aware reasoning, data scarcity, and safety-critical evaluation.
Recent frameworks such as NELV demonstrate a promising path toward fully autonomous and intelligent UAV operations by connecting natural-language tasking with end-to-end flight execution, including route planning, path generation, control, and real-time monitoring. However, extending these capabilities to large-scale UAV swarms introduces additional coordination and control challenges. Using LLMs for intelligent, scalable, and context-aware swarm coordination offers a compelling approach to addressing these challenges and advancing the next generation of autonomous UAV systems.
Table VII: Comparative overview of recent LLM-assisted UAV planning, coordination, and swarm control frameworks, highlighting their autonomy layers, application domains, core LLM functions, and demonstrated performance benefits across diverse mission scenarios.
@@ III-B LLM-Assisted UAV Swarm Control and Execution
In this context, LLMs have emerged as a powerful interface connecting human intent, high-level reasoning, and low-level swarm execution. By enabling natural language interaction, structured reasoning, and real-time code or policy generation, LLMs offer more intuitive and flexible swarm control compared to traditional interfaces. Unlike planning-oriented approaches, execution-level frameworks emphasize how swarm behaviors are generated, deployed, and coordinated in real-time, accounting for factors such as inference latency, distributed computation, and onboard resource limitations. This subsection focuses on LLM-assisted swarm control and execution mechanisms, including natural language-based formation control, hybrid LLM–RL architectures, distributed LLM inference across multiple UAVs, and low-latency control code generation. These approaches position LLMs not only as planners but also as integral components of the control and system stack, enabling scalable, responsive, and human-centric swarm operation.
@@
In addition, Flying Drones with LLM (TypeFly) [13] addresses latency by reducing the time required for LLMs to generate control code. Instead of using verbose languages such as Python, TypeFly introduces MiniSpec, a compact, token-efficient planning language designed for LLM-assisted control. Features such as skill probes, replanning, and stream interpreting – where actions are executed as code is generated – reduce response time by up to 62%. This significantly improves real-time responsiveness in time-critical UAV operations. Table VII provides a summary of these works.
Overall, these above studies demonstrate the rapid evolution of LLM-assisted UAV swarm intelligence. Each addresses a distinct challenge, including natural language interaction, collaborative reasoning, distributed inference, and low-latency control. Together, they indicate a future where UAV swarms can be commanded and coordinated through scalable, intelligent, and seamless LLM-assisted interfaces.
Table VIII: LLM-enhanced frameworks for UAV safety, network optimization, and autonomous intelligence, comparing methodologies, core LLM roles, and performance improvements in accident investigation, integrated communications, multi-objective optimization, open-world autonomy, and traffic management.
@@ III-C Leveraging LLMs for UAV Safety, Autonomy, and Network Optimization
Similarly, Li et al.[60] address multi-objective optimization in UAV networks equipped with Integrated Sensing and Communication (ISAC) capabilities. They propose an LLM-assisted decomposition-based multi-objective evolutionary algorithm (LEDMA) that balances communication and sensing performance. By decomposing complex, non-convex optimization problems into simpler subproblems and using LLMs as black-box search operators, the framework iteratively refines solutions for UAV deployment and power control. Numerical evaluations show that LEDMA achieves a superior Pareto front, faster convergence, and higher hypervolume than conventional algorithms, demonstrating the effectiveness of LLMs in solving complex wireless optimization problems.
@@
In addition, the educational domain has also benefited from LLM integration. That is, FLY-LLM Sim [39] is a modular laboratory platform designed for hands-on training with LAUS. The platform includes modules for advanced aerial navigation, where an LLM dynamically selects path-planning algorithms based on natural-language commands, as well as modules for analyzing vision-based cyberattacks on UAV perception systems. By combining the AirSim simulation environment with ChatGPT-4.0 for natural-language command processing and code generation, the platform provides a cost-effective, practical learning experience. User evaluations confirm its effectiveness, highlighting its value for mastering LLM–UAV integration. Table VIII provides a summary of these works.

### v2

@@ II-E1 Edge, On-Device, and Hybrid LLM Deployment
    Centralized Edge Inference: UAVs offload sensor data to an edge server hosting the LLM. Although communication-intensive, techniques such as model compression and token reduction can mitigate overhead. With proper optimization (e.g., quantization, pruning, parameter-efficient tuning), edge inference can support near–real-time decision-making for safety-critical tasks.
@@
    Local Inference (On-Device): Running LLMs directly on UAV hardware improves privacy, reduces latency, and enhances robustness to connectivity loss. Lightweight deployments, such as Aero-LLM [23], demonstrate the feasibility of using small models (e.g., OPT-125M/350M). However, onboard inference is constrained by limited computational resources, high energy consumption, and limited memory capacity, which often limit reasoning depth for complex tasks such as trajectory planning and resource allocation [63].
@@
    Split Inference (Model Partitioning): The model is divided across UAV and edge/cloud platforms. Early layers are executed onboard to process sensor inputs, while deeper layers are executed remotely. This reduces onboard computation and transmits compact intermediate representations.
@@
    Collaborative Device–Server Inference: Lightweight onboard models cooperate with powerful remote LLMs. Methods such as speculative decoding allow local draft generation followed by server-side verification, reducing bandwidth compared with full data offloading and improving robustness under limited connectivity [80].
Each paradigm entails trade-offs among latency, energy consumption, memory constraints, and communication overhead. The practicality of LLM deployment in UAV systems ultimately depends on hardware capability, network conditions, and mission requirements.
II-E2 Balancing Model Size, Memory, and Performance on the Edge
@@
The scalability of LLMs on edge hardware is fundamentally constrained by memory capacity, computational throughput, and energy availability, all of which are significantly more limited than in cloud environments. A theoretical upper bound on the number of model parameters (P) that can be accommodated on an edge device can be approximated as $P=\frac{M\times 8}{b}$, where $M$ denotes the available memory in bytes and $b$ represents the bit precision per parameter. In practice, however, this upper bound is substantially reduced by runtime overheads, activation storage, and intermediate buffers, thereby reducing the required capacity to roughly 1–10 billion parameters. This practical range is consistent with the scale of current state-of-the-art on-device models, such as Gemini Nano [21] and Qualcomm’s LLMs with up to 10 billion parameters [102]. Consequently, most real-world edge deployments operate in the 1–3 billion-parameter regime, whereas larger models require cloud offloading or distributed execution across multiple devices [64, 33]. Given the stringent memory, computational, and energy constraints of edge platforms, achieving real-time LLM inference on UAVs typically requires not only careful model scaling and optimization, but also high-speed, low-latency communication with edge or cloud servers.
II-E3 Structuring Knowledge for Reliable Retrieval
@@
Chunking is a fundamental component of modern RAG systems. As LLM context windows continue to grow, the primary bottleneck has shifted from generation to retrieval, making the structure and granularity of chunks central to overall system performance. Effective chunking breaks information into coherent and meaningful units that reflect how humans naturally organize and understand ideas, rather than relying on arbitrary text boundaries. This distinction is critical because the retrieval quality determines the context the model ultimately sees. Most RAG hallucinations arise from missing, fragmented, or poorly segmented information. Different chunking strategies, such as fixed-length, hierarchical, semantic, sliding-window, sentence-based, topic-boundary, and hybrid approaches, offer distinct strengths and tradeoffs, and their effectiveness depends heavily on the structure and nature of the underlying content. Ultimately, chunking is not merely a preprocessing step; it functions as an information architecture for AI, shaping how models perceive, retrieve, and reason over knowledge. Better chunking leads directly to more accurate, efficient, and reliable RAG systems.
II-E4 Quantization as an Inference Optimization Technique for Efficient LLMs
@@
Quantization is a widely used model compression technique that improves the efficiency of LLMs by reducing the numerical precision of their parameters, and, in some cases, their activations, from high-precision floating-point formats (e.g., FP32 or FP16) to lower-bit representations such as INT8, INT4, or INT3. This reduction in precision substantially decreases model size, memory footprint, computational complexity, inference latency, and energy consumption, making quantization particularly important for deploying LLMs on resource-constrained edge devices, including Raspberry Pi–class hardware. Quantization can be applied after training via Post-Training Quantization (PTQ), which converts the pretrained model weights without retraining, or during training via Quantization-Aware Training (QAT), which simulates low-precision arithmetic to better preserve accuracy. Despite these efficiency gains, quantization introduces trade-offs: aggressive reductions in bit width may lead to accuracy degradation, numerical instability, or additional dequantization overhead, particularly for tasks involving complex reasoning or arithmetic. As a result, effective quantization requires carefully balancing efficiency gains with acceptable performance loss to enable practical, sustainable edge AI deployment.
@@ III LLM for UAV Communications
Owing to their high mobility and flexible deployment, UAVs have been widely adopted in civil and commercial applications, including environmental monitoring, traffic management, parcel delivery, and precision agriculture [73, 85, 43, 62]. This section examines how LLMs enhance UAV communications across four key dimensions. It first explores LLM-assisted navigation, trajectory planning, and placement strategies for communication-aware UAV deployment. It then discusses swarm-level control and coordinated execution enabled by language-driven reasoning. Next, it analyzes how LLMs contribute to UAV safety, autonomy, and network optimization under dynamic and resource-constrained conditions. Finally, it reviews multi-LLM architectures designed to support scalable, collaborative, and communication-efficient UAV systems.
@@ III-A LLM-Assisted UAV Navigation, Planning, and Placement
Addressing this gap, Yuan et al.[111] introduce Next-Generation LLM for UAV (NELV), the first end-to-end framework that directly links natural-language instructions to fully executable UAV flight trajectories across short-, medium-, and long-range missions. NELV integrates five key components, LLM-as-Parser, route planning, path planning, control execution, and real-time monitoring, to enable complex tasks such as multi-UAV patrol, multi-point-of-interest delivery, and long-distance multi-hop relocation. Through three representative use cases, the framework demonstrates how LLMs can significantly reduce pilot workload while adhering to regulatory, airspace, and environmental constraints. Furthermore, the authors propose a five-level autonomy roadmap that outlines the progression from current language-based instruction parsing to fully autonomous LLM-as-Autopilot systems, identifying critical challenges related to regulation-aware reasoning, data scarcity, and safety-critical evaluation.
Recent frameworks such as NELV demonstrate a promising path toward fully autonomous and intelligent UAV operations by connecting natural-language tasking with end-to-end flight execution, including route planning, path generation, control, and real-time monitoring. However, extending these capabilities to large-scale UAV swarms introduces additional challenges for coordination and control. Using LLMs for intelligent, scalable, and context-aware swarm coordination offers a compelling approach to addressing these challenges and advancing the next generation of autonomous UAV systems.
Table V: Comparative overview of recent LLM-assisted UAV planning, coordination, and swarm control frameworks, highlighting their autonomy layers, application domains, core LLM functions, and demonstrated performance benefits across diverse mission scenarios.
@@ III-B LLM-Assisted UAV Swarm Control and Execution
In this context, LLMs have emerged as a powerful interface connecting human intent, high-level reasoning, and low-level swarm execution. By enabling natural-language interaction, structured reasoning, and real-time code or policy generation, LLMs offer more intuitive and flexible swarm control than traditional interfaces. Unlike planning-oriented approaches, execution-level frameworks emphasize how swarm behaviors are generated, deployed, and coordinated in real-time, accounting for factors such as inference latency, distributed computation, and onboard resource limitations. This subsection focuses on LLM-assisted swarm control and execution mechanisms, including natural-language-based formation control, hybrid LLM–RL architectures, distributed LLM inference across multiple UAVs, and low-latency code generation for control. These approaches position LLMs not only as planners but also as integral components of the control and system stack, enabling scalable, responsive, and human-centric swarm operation.
@@
In addition, Flying Drones with LLM (TypeFly) [14] addresses latency by reducing the time required for LLMs to generate control code. Instead of using verbose languages such as Python, TypeFly introduces MiniSpec, a compact, token-efficient planning language designed for LLM-assisted control. Features such as skill probes, replanning, and stream interpretation – in which actions are executed as code is generated – reduce response time by up to 62%. This significantly improves real-time responsiveness in time-critical UAV operations. Table V provides a summary of these works.
Overall, the above studies demonstrate the rapid evolution of LLM-assisted UAV swarm intelligence. Each addresses a distinct challenge, including natural language interaction, collaborative reasoning, distributed inference, and low-latency control. Together, they indicate a future in which UAV swarms can be commanded and coordinated via scalable, intelligent, and seamless LLM-assisted interfaces.
Table VI: LLM-enhanced frameworks for UAV safety, network optimization, and autonomous intelligence, comparing methodologies, core LLM roles, and performance improvements in accident investigation, integrated communications, multi-objective optimization, open-world autonomy, and traffic management.
@@ III-C Leveraging LLMs for UAV Safety, Autonomy, and Network Optimization
Similarly, Li et al.[55] address multi-objective optimization in UAV networks equipped with Integrated Sensing and Communication (ISAC) capabilities. They propose an LLM-assisted decomposition-based multi-objective evolutionary algorithm (LEDMA) that balances communication and sensing performance. By decomposing complex, non-convex optimization problems into simpler subproblems and using LLMs as black-box search operators, the framework iteratively refines solutions to UAV deployment and power control problems. Numerical evaluations show that LEDMA achieves a superior Pareto front, faster convergence, and higher hypervolume than conventional algorithms, demonstrating the effectiveness of LLMs in solving complex wireless optimization problems.
@@
In addition, the educational domain has also benefited from LLM integration. That is, FLY-LLM Sim [36] is a modular laboratory platform designed for hands-on training with LAUS. The platform includes modules for advanced aerial navigation, where an LLM dynamically selects path-planning algorithms based on natural-language commands, as well as modules for analyzing vision-based cyberattacks on UAV perception systems. By combining the AirSim simulation environment with ChatGPT-4.0 for natural-language command processing and code generation, the platform provides a cost-effective, practical learning experience. User evaluations confirm its effectiveness and highlight its value for mastering LLM–UAV integration. Table VI provides a summary of these works.